A particle is thrown with a velocity u at an angle theta from the horizontal. Another particle is thrown with the same velocity at an angle a from the vertical. The ratio of times of flight of the two particles will be

Grade:

1 Answers

Eshan
5 years ago
Dear studnet,

The time of flight of a projectile thrown at an angle\thetafrom horizontal is given as
t=\dfrac{2usin\theta}{g}
For projectile thrown forming same angle with vertical, the angle that it foms with horizontal is90^{\circ}-\theta
Hence the time of flight ist'=\dfrac{2usin(90^{\circ}-\theta)}{g}=\dfrac{2ucos\theta}{g}
Hence\dfrac{t}{t'}=tan\theta